OpenStudy (anonymous):

Please Help!!!! Fan and Medal!!!!!!!!!! Solve, using linear combination. 3x – y = 4 2x + y = 6 A. (3, 0) B. (2, 2) C. (0, –4) D. (–1, –7)

OpenStudy (anonymous):

@mathstudent55 , @SolomonZelman , @joshhfers , @paki , @AccessDenied

OpenStudy (arbershabani97):

just plug them in and see which one works ( for example A) 3,0 3*3-0=4 -----not correct

OpenStudy (anonymous):

ok

OpenStudy (accessdenied):

That is one way to do it, and always works out by trial and error. To use linear combinations, we are trying to eliminate a variable by adding the equations together where one variable is added and subtracted equally. So in our case, equation 1 has a -y and equation 2 has a + y. If we added the left sides of the equations, those will cancel out, right? (3x - y ) + (2x + y) = 4 + 6 + y and - y cancel, 2x and 3x are like terms that add together... 5x = 10 This allows you to solve for one variable, and then substitute your x-value to find the y-value in either equation.